Find the length of the hypotenuse of a right triangle with legs of lengths 5 and 12. If necessary, round your answer to two decimal places.
Answer by stanbon(75887)

Find the length of the hypotenuse of a right triangle with legs of lengths 5 and 12
---------------------
h^2 = 5^2 + 12^2
h = sqrt(25+144)
h = sqrt(169)
h = 13
=================
Cheers,
Stan H.